OpenStudy (anonymous):

Evaluate the indefinite integral of e^4x

OpenStudy (zzr0ck3r):

(1/4)e^(4x)<--- what would the derivative of this be?

OpenStudy (zzr0ck3r):

@danielashvartz ?

OpenStudy (lgbasallote):

\[\large \int e^{4x}dx\] let u = 4x so du = 4dx so du/4 = dx now substitute.. \[\implies \int e^u \frac{du}{4}\] \[\implies \frac 14 \int e^u du\] does that help?

OpenStudy (zzr0ck3r):

I must have scared him away:P

OpenStudy (unklerhaukus):

\[\int e^{4x}=\frac {e^{4x}}4+c\]

OpenStudy (lgbasallote):

...yes @UnkleRhaukus ....way to go....